Two metal balls are hung from strings.  The first ball, 0.13 kg, is pulled to the side and released so that it strikes the second ball with a speed of 1.66 m/s.  The two balls stick together and the combined mass has a speed of 0.26 m/s after collision.  What is the mass of the second ball?
Answer #1

initial momentum = m1 v1 + m2*v2 = 0.13 * 1.66 + m2*0 = 0.216 kgm/s

final momentum = (m+m2)*v = (0.13+m) * 0.26 kgm/s

Conservation of momentum

initial momentum = final momentum

so, 0.216 = (0.13+m)*0.26

so 0.13+m= 0.83

so m = 0.7 kg -->Answer
Answer #2

SOLUTION 


Momentum before strike = Momentum after strike

=> m1 v1 + m2 v2 = (m 1 + m2) v

=> (0.13 * 1.66) + m2 * 0 = (0.13 + m2) * 0.26

=> m2 = (0.13 * 1.66) / 0.26  - 0.13

=> m2 = mass of second ball = 0.7 kg (ANSWER).

answered by: Tulsiram Garg